OpenStudy (anonymous):

determine the number and type of solutions for x^2 + 9x +7=0

OpenStudy (anonymous):

well it's a quadratic so there will be two solutions. You can tell because of the \[x^2\]

OpenStudy (anonymous):

2 irrational, rational, or complex?

OpenStudy (anonymous):

You can use the quadratic formula to solve for x \begin{array}{*{20}c} {x = \frac{{ - b \pm \sqrt {b^2 - 4ac} }}{{2a}}} & {{\rm{when}}} & {ax^2 + bx + c = 0} \\ \end{array}

OpenStudy (anonymous):

Hmm well you can try solving it see what type of values you will get if it's hard to tell by looking at it.

OpenStudy (cwrw238):

discriminant = 81 - 28 = 53 so thers 2 rational roots

OpenStudy (anonymous):

thank you!

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

How did you find the discriminant?

OpenStudy (cwrw238):

b^2 - 4ac here b = 9,a=1 and c = 7 so its 9^2 - 4*1* 7

OpenStudy (anonymous):

that answer was wrong..btw o-o

OpenStudy (anonymous):

Right if the discriminant is higher than zero then we know that it can't be a complex number. Btw if the answer is wrong there is only one other possible solution.

OpenStudy (cwrw238):

if D > 0 then there are 2 roots - oh right they are not necessarily rational

OpenStudy (cwrw238):

in fact since D = sqrt53 there are both irrational

OpenStudy (cwrw238):

i should have said ( 2 posts ago) D>0 means there are 2 REAL roots which could be rational or irrational